OpenStudy (anonymous):

Find y'. y=2√x- 1/2√x

OpenStudy (anonymous):

\[d/dx(x ^{n})=nx ^{n-1}\]

OpenStudy (anonymous):

\[\sqrt{x}=x ^{1/2}\]

OpenStudy (anonymous):

so now i think u can find the derivative of \[x ^{1/2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so would the answer be \[ \frac{ 1 }{ \sqrt{x}\ } + \frac{ 1 }{ 4x \sqrt{x} }\]